# Grokking Group Multiplication with Cosets

### v1

@@ Abstract
We use the group Fourier transform over the symmetric group $S_{n}$ to reverse engineer a 1-layer feedforward network that has ”grokked” the multiplication of $S_{5}$ and $S_{6}$. Each model discovers the true subgroup structure of the full group and converges on circuits that decompose the group multiplication into the multiplication of the group’s conjugate subgroups. We demonstrate the value of using the symmetries of the data and models to understand their mechanisms and hold up the “coset circuit” that the model uses as a fascinating example of the way neural networks implement computations. We also draw attention to current challenges in conducting mechanistic interpretability research by comparing our work to Chughtai et al. [6] which alleges to find a different algorithm for this same problem.
Machine Learning, Grokking, Mechanistic Interpretability, Group Theory, Harmonic Analysis
@@ 1 Introduction
The program of modern mechanistic interpretability was first clearly articulated in [44]: deep neural networks use directions in neuron activation space as features and perform tasks by combining those features in neural “circuits.” The features that neural networks learn and the circuits that they form do not have to be opaque. With care and effort researchers can peer into the black box of deep models and understand how they accomplish the tasks that they are trained to accomplish.
We follow prior work in circuit-level mechanistic interpretability and study a model and a task that we can understand completely [6, 23, 40, 41]. In so doing we hope to build an understanding of the the types of features that neural networks form and the precise mechanisms with which neural networks combine and transform those features. We hope that by reverse-engineering the function of each neuron in a model that performs a simple task we will contribute to a foundation of understanding from which to study the mechanisms of massive models.
Figure 1: The coset circuit that the model uses to perfectly generalize the group multiplication. The circuit stores information about coset membership for the left and right permutations in the embeddings. The linear layer queries the embeddings and combines the information to determine whether or not the product is in a specific coset. Coset membership is encoded in the sign of the activations.
In this paper we return to the subject of Chughtai et al. [6] and analyze one hidden layer fully connected networks of varying widths that have perfectly learned the task of group multiplication. We found their analysis incomplete, however, and build on it by using the full Group Fourier transform to completely reverse-engineer the most common mechanism converged networks use to implement group multiplication. Though there are in fact many different neural circuits that the networks form, we can describe the vast majority using the cosets of $S_{n}$ that we refer to as the coset circuit. Cosets are subsets of a group that have specific properties in relation to each other that the model can use to perfectly generalize in the task of group multiplication (see Section 3.1 for concrete examples of cosets for permutation groups or Appendix D for a formal definition).
The high level steps of the coset circuit, given two permutations $\sigma_{l},\;\sigma_{r}$, are:
  1. 1.
    In the left and right embedding layers, the model stores information about membership in dozens of left and right cosets.
  2. 2.
    In the linear layer, the embeddings are queried to get the coset membership for $\sigma_{l}$ and $\sigma_{r}$. The linear layer combines this information to get the probable coset membership of their product $\sigma_{l}\sigma_{r}$.
  3. 3.
    In the unembedding layer, the identity of $\sigma_{l}\sigma_{r}$ is determined as the only permutation in $S_{n}$ that satisfies all of the different coset memberships computed by the linear layer in step 2.
1.1 Our Contributions
@@
  1. 1.
    We completely reverse engineer a 1-hidden layer fully-connected network trained on the permutation groups $S_{5}$ and $S_{6}$.
  2. 2.
    We present a novel application of the Group Fourier Transform to circuit-level mechanistic interpretability.
  3. 3.
    We apply a methodology inspired by Geiger et al. [21] to use causal interventions thoroughly test all of the properties of our proposed circuit.
  4. 4.
@@ 2 Related Work
Mechanistic Interpretability We analyze the mechanisms of neural networks at the circuit level of individual and small groups of neurons, drawing directly on the work of [40, 41, 45, 59]. Our work builds directly on “A Toy Model of Universality” by Chughtai et al. [6]. We recreated precisely their experimental setup for the groups $S_{5}$ and $S_{6}$, but found their analysis incomplete and could not find a mechanism implementing their proposed “Group Composition via Representations” algorithm. For a thorough comparison of our work and [6], see Appendix F.
Grokking Our choice of studying models trained on small permutation groups was made in no small part because of the appearance of “grokking,” wherein the model first memorizes a training set and then much later generalizes to the held out data perfectly. Grokking was first identified by Power et al. [47] and has been well studied for its counter-intuitive training dynamics [31, 38, 52, 49, 57, 56]. We conducted all the analysis on fully grokked models with perfect test accuracy.
Group Theory Our analysis owes much to the seminal monograph “Group Representations in Probability and Statistics” [12] and the varied literature building on it [8, 30, 29, 26, 28]. We implement in Pytorch the Fast Fourier transform algorithm over the symmetric group $S_{n}$ developed by Clausen & Baum [7]. Our implementation is heavily inspired by the example in Julia of Plumb et al. [46].
3 Background
@@ 3.1 Permutations and the Symmetric Group
In this section we will give a concrete exposition of some concepts of group theory as they are concretely realized in the permutation groups that we study. For a more formal introduction to group theory, please refer to Appendix D.
A permutation of $n$ elements is a map $\sigma$ that sends one ordering of the $n$ elements to a different ordering, for example the order-reversing permutation on four elements would be:
@@
and the identity permutation, denoted $e$, changes nothing:
@@
We write specific permutations by identifying them with the image of their action on the elements $\{1,2,\dots,n\}$ in their standard order. For the above example, we would simply denote the order reversing permutation on four elements as (4 3 2 1)4321(4\;3\;2\;1)( 4 3 2 1 ).
@@
What we see is that first applying $\tau$ and then $\sigma$ has the same effect as just applying the permutation (4 1 2 3)4123(4\;1\;2\;3)( 4 1 2 3 ). Additionally every permutation $\sigma$ has an inverse $\sigma^{-1}$ such that $\sigma\sigma^{-1}=e$. These properties makes all of the permutations on $n$ elements a group called the symmetric group, which we write $S_{n}$.
There are six permutations in $S_{4}$ that ”fix” $4$:
@@
These six permutations form a subgroup of $S_{4}$ because multiplying any two permutations that fix four results in another permutation that fixes four. You can see that these six permutations are homomorphic to $S_{3}$ by simply “forgetting” about the $4$ that is fixed in the fourth position.
@@
This coset is characterized by every element having $4$ in the first position, because $\tau$ switches positions the first and fourth positions and $S_{3}$ always has $4$ in the fourth position. We would get a coset with all of the elements of $S_{4}$ with $4$ in the third position if we multiplied $S_{3}$ the left by the transposition that switches three and four.
There are also right cosets where every element in a subgroup is multiplied from the right. The elements of $S_{3}\tau$ are:
@@
This right coset is characterized by every element having $1$ in the fourth position, because $S_{3}$ fixes the element in the fourth position and $\tau$ has $1$ in the fourth position.
@@
If a subgroup is invariant to conjugation then it is a normal subgroup. The only normal subgroup of $S_{n}$ for $n>4$ is the alternating group $A_{n}$ of even permutations.
A transposition is a permutation $\tau\in S_{n}$ that switches (“transposes”) two elements $i,j\in[n]$ and leaves the remaining elements fixed. Every element of $S_{n}$ can be decomposed into a product of transpositions. For a given permutation this decomposition is not unique, but the number of transpositions in the decomposition is constant. The permutations that have an even number of transpositions are referred to as “even” permutations and those with an odd number are “odd.” $A_{n}$ is made up of the permutations that have an even number of transpositions. For normal subgroups the left and right subgroups are equal, so $\sigma A_{n}=A_{n}\sigma$.
@@
Figure 2: Model Architecture: we follow the model architecture used by [6]. The one-hot vectors of left and right permutations pass through separate embeddings. We concatenate the embeddings and pass them through a single fully-connected hidden layer with ReLU activations. An unembedding matrix transforms the activations into logits.
@@ 3.2 Fourier Transform over Groups
The Discrete Fourier Transform (DFT) converts a function $f$ defined on $\{0,1,\dots,n-1\}$ to a complex-valued function via the formula:
@@
This basis has a number of useful properties, that stem from how well-fit it is to the structure of the domain. There are many functions that appear complicated in the time domain that are relatively simple or sparse in the frequency domain.
@@
The interpretation of the DFT as being over the cyclic groups can be generalized to non-commutative groups. We go over the construction in Appendix D.6 and E. The high level interpretation, however, is the same. For functions from $S_{n}\rightarrow\mathbb{C}$ there is an orthonormal basis that is equivariant to translations and convolutions. The frequencies for the Fourier transform over $S_{n}$ are given by the partitions of $n$. The “highest” frequencies can be interpreted as representing functions that are constant on permutations that all agree on a small number of elements of $[n]$ [16].
@@ 4 Model Architecture
We study a fully connected feed-forward network that is trained to learn the multiplication of finite permutation groups. As shown in Figure 2, the model contains separate left and right embeddings, followed by a fully connected linear layer with ReLU activations, and an unembedding layer.
@@
  • Two embedding matrices, $\mathbf{E}_{l},\;\mathbf{E}_{r}$ with dimensions $(d,\;|G|)$, where $d$ is embedding dimension.
@@ 5 Coset Circuits
In this section we will describe the precise mechanism of the coset circuits that the models use to implement multiplication in $S_{5}$ and $S_{6}$.
First in Section 5.1 we will describe how in the linear layer the simplest circuit, the ”Sign Circuit,” implements the multiplication of the two cosets of the alternating group $A_{n}$.
In Section 5.2 we will describe how the linear layer uses coset information with conjugate subgroups such as $S_{4}<S_{5}$.
And in Section C.3 we discuss the distribution of different circuits and subgroups that the model converges on over many different runs.
5.1 Sign Circuit
(a) Weight Analysis in Fourier Basis (Left)
(b) Weight Analysis in Fourier Basis (Right)
(c) Activations of all permutations combinations pre-relu
(d) Activations of all permutations combinations post-relu
Figure 3: Four subfigures in two subgroups
The even permutations form a subgroup called the alternating group $A_{n}$. $A_{n}$ has two cosets: the group itself $A_{n}$ and all of the odd permutations $\tau A_{n}$. The sign map $\operatorname{sgn}$ is given by:
@@ 5.1 Sign Circuit
The sign map respects multiplication in $S_{n}$. For any $\sigma,\rho\in S_{n}$, the sign of their product is the product of their signs: $\operatorname{sgn}(\sigma\rho)=\operatorname{sgn}(\sigma)\operatorname{sgn}(\rho)$.
There are three steps to each neuron’s mechanism. In what follows, $x$ is the magnitude of the pre-activation, $\sigma$ refers to a permutation, and the subscripts $l,r,t$ on $\sigma$ either refer to the left input permutation, the right input permutation, or the target permutation $\sigma_{t}=\sigma_{l}\sigma_{r}$, respectively.
  1. 1.
    The neuron weights corresponding to the left and right permutations “read off” the pre-activations from the embeddings. The pre-activations for the left and right take on one of two values: $x\operatorname{sgn}(\sigma)$ or $-x\operatorname{sgn}(\sigma)$.
  2. 2.
    The left and right pre-activations are added together through linear layer, so that the full pre-activations now take on one of three values: $-2x$, 00, or $2x$.
  3. 3.
    The ReLU activation function cuts off the pre-activations if they are $-2x$ and otherwise passes them along unchanged to the unembedding layer.
Let us take as a concrete example the neuron shown in Fig. 3, where the left pre-activations $L(\sigma)=-\operatorname{sgn}(\sigma)x$ and the right pre-activations $R(\sigma)=\operatorname{sgn}(\sigma)x$. The full action of the neuron is given by $\max(0,L(\sigma_{l})+R(\sigma_{r})$ and there are three cases:
@@
    $\operatorname{sgn}(\sigma_{l})=-1,\;\operatorname{sgn}(\sigma_{r})=1\Rightarrow\operatorname{sgn}(\sigma_{l}\sigma_{r})=-1$. In this case $L(\sigma_{l})$ and $R(\sigma_{r})$ reinforce each other and sum to a positive value. Since $02x>$ 0, the activation value is $2x$.
@@
    $\operatorname{sgn}(\sigma_{l})=1,\;\operatorname{sgn}(\sigma_{r})=-1\Rightarrow\operatorname{sgn}(\sigma_{l}\sigma_{r})=-1$. Like in (2) the product $\sigma_{l}\sigma_{r}$ is an odd permutation and $L(\sigma_{l})$ and $R(\sigma_{r})$ constructively interfere, though this time $L(\sigma_{l})+R(\sigma_{r})=-2x<0$. Thus ReLU clips the pre-activation and sends it to 00, so despite the information that $\sigma_{l}\sigma_{r}$ is even is present in the pre-activations, the activation that gets passed to the unembedding layer is identical to an even permutation as in (1).

### v2

@@ Abstract
The complex and unpredictable nature of deep neural networks prevents their safe use in many high-stakes applications. There have been many techniques developed to interpret deep neural networks, but all have substantial limitations. Algorithmic tasks have proven to be a fruitful test ground for interpreting a neural network end-to-end. Building on previous work, we completely reverse engineer fully connected one-hidden layer networks that have “grokked” the arithmetic of the permutation groups $S_{5}$ and $S_{6}$. The models discover the true subgroup structure of the full group and converge on neural circuits that decompose the group arithmetic using the permutation group’s subgroups. We relate how we reverse engineered the model’s mechanisms and confirmed our theory was a faithful description of the circuit’s functionality. We also draw attention to current challenges in conducting interpretability research by comparing our work to Chughtai et al. [4] which alleges to find a different algorithm for this same problem.
Machine Learning, Grokking, Interpretability, Group Theory, Harmonic Analysis
@@ 1 Introduction
Many methods have been proposed to render deep neural networks interpretable. There is both an academic interest in understanding how neural networks do what they do and a societal interest in ensuring that decisions made by such models are sound, unbiased, and subject to human review. These concerns are not new, nor are they unique to deep neural networks. Many of the techniques developed (such as SHAP values [33], saliency maps [52], gradient attribution [51], dimension reduction [61], etc…) are still widely used today, but there is an understanding that such methods must be used as just one part of a careful analysis. Naive applications of even the most sophisticated algorithms will give misleading results [1, 2, 9, 25].
Mechanistic interpretability seeks to find “neural circuits” within deep neural networks, small sub-networks that act as connected computation graphs and accomplish a task. In “toy” (highly constrained) settings mechanistic interpretability has been successful, with multiple examples where the inner workings of neural networks have been successfully reverse engineered end-to-end [20, 40, 41, 49, 63]. There have also been encouraging early successes in finding interpretable circuits within real-world models [18, 32, 35, 44, 57], but there is already work emerging that illustrates how neural networks can resist common “mechanistic interpretability” methods [14, 34, 60].
The toy interpretability projects that have succeeded have done so in large part because a distinct ground truth circuit that encodes the true nature of the task or environment emerged in the model. We build on this tradition and study a model that has perfectly learned to multiply permutations of five and six elements, which in mathematics is known as the symmetric groups $S_{5}$ and $S_{6}$, which are deeply studied and well-understood objects [8, 10, 15]. We succeed in completely reverse engineering the model and enumerating the diverse circuits that it converges on to implement the multiplication of the symmetric group. Our work does not, however, represent an unmitigated success for the project of mechanistic interpretability. The prior work of Chughtai et al. [4] studied the exact same model and setting, but came to completely different conclusions. Understanding why our and Chughtai et al. [4]’s interpretations of the same data diverged required extensive effort (see Appendix 7 for a thorough comparison). We find that even in a setting as simple and well understood as group arithmetic, it is incredibly difficult to do interpretability research and be confident about one’s conclusions.
@@
  • We completely reverse engineer a one-hidden layer fully-connected network trained on the permutation groups $S_{5}$ and $S_{6}$.
  • We apply a methodology inspired by Geiger et al. [17] to use causal experiments to thoroughly test all of the properties of our proposed circuit.
@@ 2 Related Work
Mechanistic Interpretability Interpreting and reverse engineering the mechanism used to complete a given task is an active field in interpretability. Analysis of such mechanisms and circuits are discovered mainly through a top-down approach of causal mediation analysis. In the previous work Hanna et al. [21], Meng et al. [36], Tigges et al. [54], Wang et al. [58], the circuits are composed at the “component level” using the feed-forward layer and attention heads. We analyze the mechanisms of neural networks at the circuit level of individual and small groups of neurons, drawing directly on the work of Nanda et al. [40, 41], Olah et al. [43], Quirke & Barez [49], Zhong et al. [64], Zhang et al. [63]. Our work builds directly on “A Toy Model of Universality” by Chughtai et al. [4]. We recreated precisely their experimental setup for the groups $S_{5}$ and $S_{6}$, though we came to different conclusions.
Grokking The models we study exhibit “grokking”, wherein the model first memorizes the training set and then much later generalizes to the held out data perfectly. Grokking was first identified by Power et al. [47] and has been well studied for its counter-intuitive training dynamics [31, 37, 55, 50, 62, 59]. We conducted all the analysis on fully grokked models with perfect test accuracy, as models that show this behavior have often formed clean generalizing circuits that are more easily interpreted [20, 40].
Group Theory We used many of the tools of group theory for our analysis, in particular the well-developed representation theory of the symmetric group. Tools for analyzing data on groups are well-laid out in Clausen & Baum [5], Cohen & Welling [6], Diaconis [8], Kondor [29], Kondor & Trivedi [30], Huang et al. [24], Karjol et al. [27], Plumb et al. [46].
3 Mathematical Preliminaries
Figure 1: Model Architecture: we follow the model architecture used by Chughtai et al. [4]. The one-hot vectors of left and right permutations pass through separate embeddings. We concatenate the embeddings and pass them through a single fully-connected hidden layer with ReLU activations. An unembedding matrix transforms the activations into logits.
This paper requires a familiarity with functions on groups, a topic that is uncommon in machine learning research. In this section we give an overview of the major concepts as they are realized in the permutation groups that we study. For a more formal introduction to group theory, please refer to Appendix D.
@@ 3.1 Permutations and the Symmetric Group
A permutation of $n$ elements is a map $\sigma$ that sends one ordering of $n$ elements to a different ordering. For example the order-reversing permutation on four elements would be:
@@
The identity permutation, denoted $e$, leaves the ordering unchanged:
@@
We refer to specific permutations by identifying them with the image of their action on the elements $[n]\coloneqq\{1,2,\dots,n\}$ in increasing order. For the above example we would simply denote the order reversing permutation on four elements as (4 3 2 1)4321(4\;3\;2\;1)( 4 3 2 1 ).
@@
First applying $\tau$ and then $\sigma$ has the same effect as just applying the permutation (4 1 2 3)4123(4\;1\;2\;3)( 4 1 2 3 ). Additionally every permutation $\sigma$ has an inverse $\sigma^{-1}$ such that $\sigma\sigma^{-1}=e$. These properties makes all of the permutations on $n$ elements a group called the symmetric group, which we write $S_{n}$.
There are six permutations in $S_{4}$ that do not change the position of $4$:
@@
These six permutations form a subgroup of $S_{4}$ because multiplication is closed within that subset, multiplying any two permutations that leave $4$ unchanged results in another permutation that leaves $4$ unchanged. You can see that these six permutations are isomorphic to $S_{3}$ by simply “forgetting” about the $4$ that is fixed in the fourth position. In the paper, we will refer to the subgroup of $S_{n}$ isomorphic to $S_{n-1}$ that leaves element $i$ fixed as $H_{i}$.
One of the simplest types of permutations is a “transposition,” a permutation $\tau\in S_{n}$ that switches (“transposes”) two elements $i,j\in[n]$ and leaves the remaining elements fixed. Every element of $S_{n}$ can be decomposed into a product of transpositions. A given decomposition of a permutation is not unique, but the number of transpositions in the decomposition is an invariant of the permutation. For a permutation $g\in S_{n}$ if a set of transpositions $\tau_{1}\tau_{2}\dots\tau_{k}=g$, then every possible such set of transpositions will also have $k$ elements. The permutations that have an even number of transpositions are referred to as “even” permutations and those with an odd number are “odd.” The set of all even permutations in $S_{n}$ is a subgroup referred to as the “alternating group” $A_{n}$.
@@
This coset is characterized by every element having $4$ in the first position. Every element of $H_{4}$ has $4$ in the fourth position and $\tau$ switches the first and fourth positions. For any $h\in H_{4}$, $h\tau$ has $4$ in the first position because $\tau$ moves it from the fourth. We would get a coset with all of the elements of $S_{4}$ with $4$ in the third position if we multiplied $H_{4}$ on the left by the any permutation that switches three and four.
There are also right cosets where every element in a subgroup is multiplied from the right. The elements of $H_{4}\tau$ are:
@@
This right coset is characterized by every element having $1$ in the fourth position.
@@
If a subgroup is invariant to conjugation it is a normal subgroup. The only normal subgroup of $S_{n}$ for $n>4$ is the alternating group $A_{n}$ of even permutations.
@@ 3.2 Fourier Transform over Groups
Though Group Fourier Transform is not central to our presentation of the coset circuit, it was an important tool that we used to analyze the the activations of the trained models. It is also a critical part of [4]. We introduce the concepts here and go over the the similarities and differences between our work and [4] in Section 7.
We begin with a presentation of the Discrete Fourier Transform (DFT), and then present the Group Fourier Transform by analogy. The DFT converts a function $f$ defined on $\{0,1,\dots,n-1\}$ to a complex-valued function via the formula:
@@
The interpretation of the DFT as being over the cyclic groups can be generalized to non-commutative groups. We go over the construction in Appendix E and F. The high level interpretation, however, is the same. For functions from $S_{n}\rightarrow\mathbb{C}$ there is an orthonormal basis that is equivariant to translations and convolutions. The frequencies for the Fourier transform over $S_{n}$ are given by the partitions of $n$. The “highest” frequencies can be interpreted as representing functions that are constant on permutations that all agree on a small number of elements of $[n]$ [12].
@@ 4 Model Architecture
As shown in Figure 1, the model we study contains separate left and right embeddings, followed by a fully connected linear layer with ReLU activations, and an unembedding layer. We use the same architecture as in [4] to enable consistent comparisons.  [^1] All code necessary for reproducing results and analysis is available at https://www.github.com/dashstander/sn-grok
@@
  • Two embedding matrices, $\mathbf{E}_{l},\;\mathbf{E}_{r}$ with dimensions $(d,\;|G|)$, where $d$ is embedding dimension. $S_{n}$ is non-abelian, i.e. not commutative, and the separate embeddings are to give the model extra capacity.
@@
Throughout the paper will refer to the values $\mathbf{L}\mathbf{E}_{l}\mathbf{x}_{g}$, $\mathbf{R}\mathbf{E}_{r}\mathbf{x}_{h}$, and their sum as “pre-activations” to denote that the ReLU activation function has not been applied. Post-ReLU values we refer to as “activations.”
@@ 5 Coset Circuits
5.1 Sign Neurons Implement the Sign Circuit
Figure 2: A diagram showing the four possible paths through a single neuron (i.e. one row of $\mathbf{R}\mathbf{E}_{r}$) that implements part of a “sign circuit.” The model stores whether a permutation is “even” or “odd” in the embeddings, represented in the left or right pre-activation values. The pre-activations are added together and then the ReLU activation is applied. The neuron only fires when the left permutation is even and the right is odd. If the neuron does not fire, then in $1/3$ cases the product is odd and $2/3$ it is even.
The even permutations form a subgroup called the alternating group $A_{n}$. The two cosets of $A_{n}$ are the group itself and all of the odd permutations, $\tau A_{n}$. The multiplication of even and odd permutations has similar features to the addition of even and odd integers (hence the name). The sign map on a permutation in $S_{n}$, $\operatorname{sgn}$, is given by:
@@ 5.1 Sign Neurons Implement the Sign Circuit
An “even” permutation that is in $A_{n}$ is mapped to $1$ and an “odd” permutation not in $A_{n}$ is mapped to $-1$. For any $\sigma,\rho\in S_{n}$, the sign of their product is the product of their signs: $\operatorname{sgn}(\sigma\rho)=\operatorname{sgn}(\sigma)\operatorname{sgn}(\rho)$.
The one-layer model that we train uses this relationship to help solve the general group multiplication. Every single model we trained had at least two neurons dedicated to encoding the sign of the permutation product. Though the model cannot use the alternating group to completely solve multiplication in $S_{n}$, this sign circuit is emblematic of the general coset circuits the model forms.
Consider the neuron shown in Fig. 2. The left pre-activations are given by $L(\sigma)=\operatorname{sgn}(\sigma)$ and the right pre-activations are $R(\sigma)=-\operatorname{sgn}(\sigma)$. The full action of the neuron is given by $\texttt{ReLU}(L(\sigma_{l})+R(\sigma_{r}))$ and there are three cases:
@@
    $\operatorname{sgn}(\sigma_{l})=-1,\;\operatorname{sgn}(\sigma_{r})=1\Rightarrow\operatorname{sgn}(\sigma_{l}\sigma_{r})=-1$. In this case $L(\sigma_{l})$ and $R(\sigma_{r})$ reinforce each other and sum to a positive value. Since $2>0$, the activation value is $2$.
@@
    $\operatorname{sgn}(\sigma_{l})=1,\;\operatorname{sgn}(\sigma_{r})=-1\Rightarrow\operatorname{sgn}(\sigma_{l}\sigma_{r})=-1$. Like in (2) the product $\sigma_{l}\sigma_{r}$ is an odd permutation and $L(\sigma_{l})$ and $R(\sigma_{r})$ constructively interfere, though this time $L(\sigma_{l})+R(\sigma_{r})=-2$, which is less than 00. Thus ReLU clips the pre-activation and sends it to 00.